Numerical Analysis I, Fall 2011 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 05

Assigned Oct 14, 2011.

Correction for section 4.3, prob 18: The coefficients should be the same as problem 17.

1. Section 4.4: Problems 14, 16, 17, 19.

2. Section 4.7: Problems 1(f), 2(f), 5, 7, 8.